OpenStudy (anonymous):

Jane has $5000. She invests a portion of her money at a simple interest rate of 5% and the rest of her money at a simple interest rate of 6.2%. After one year, the total interest earned on these investments is $292.00. How much money did she invest at 5%?

OpenStudy (anonymous):

Suppose the invested amounts are x and y, at five and six percent, respectively. We know\[x+y=5000 \implies y=5000-x\]That allows us to say\[.05x+.06y=292 \implies 5x+6(5000-x)=29200 \implies 800=x\]It seems the amount invested at five percent is about eight hundred dollars.